OpenStudy (anonymous):

what is the solution of this equation? -7 + 1/2z = -4 explain.

OpenStudy (anonymous):

its one half not a division sign

OpenStudy (anonymous):

Haha ok :)

OpenStudy (anonymous):

Let's see, so -7+1/2z=-4 Add 7 to both sides, first off, so that we can get the variable on one side.

OpenStudy (anonymous):

So you have 1/2z=3

OpenStudy (anonymous):

*would

OpenStudy (anonymous):

hold on i need paper

OpenStudy (anonymous):

Ok :P

OpenStudy (anonymous):

1/2z = -4 +7 1/2z = 3 z = 2 * 3 z = 6

OpenStudy (anonymous):

how do i know that 1/2 = 3???

OpenStudy (anonymous):

Well, when you added 7 to each side you got 1/2z=3 Then you would need to dived each side by the reciprocal of 1/2, which is 2 So you would have z=2(3) Z=6 :) Does that make sense?

OpenStudy (anonymous):

ohhh well thanks, and can you help me with some other questions? ill close this one though

OpenStudy (anonymous):

Sure :)